At a total cost of $6,030,000, Herrera Corporation acquired 224,000 shares of Tran Corp. common stock as a long-term investment. Herrera Corporation uses the equity method of accounting for this investment. Tran Corp. has 800,000 shares of common stock outstanding, including the shares acquired by Herrera Corporation.

Required:

A. Journalize the entries by Herrera Corporation on December 31 to record the following information (refer to the Chart of Accounts for exact wording of account titles):
1. Tran Corp. reports net income of $933,000 for the current period.
2. A cash dividend of $0.27 per common share is paid by Tran Corp. during the current period.
B.

Why is the equity method appropriate for the Tran Corp. investment?

A. Journalize the entries by Herrera Corporation on December 31 to record the following information. Refer to the Chart of Accounts for exact wording of account titles.

PAGE 10

JOURNAL

ACCOUNTING EQUATION

DATE DESCRIPTION POST. REF. DEBIT CREDIT ASSETS LIABILITIES EQUITY

1

2

3

4

B. Why is the equity method appropriate for the Tran Corp. investment?

Herrera’s investment in Tran Corp. represents of the outstanding shares of Tran Corp. An investment amount between     and     of the outstanding common stock of the investee is presumed to represent significant influence. The equity method is when the investor can exercise significant influence over the investee.
Answer #1
% of ownership 224000/800000 = 28%
A
Date Description Post ref. Debit Credit Assets Liabilities Equity
December 31 Investment in Tran Corp. stock 261240 =933000*28%
      Income of Tran Corp. 261240
December 31 Cash 60480 =224000*0.27
      Investment in Tran Corp. stock 60480
B
Herrera’s investment in Tran Corp. represents 28% of the outstanding shares of Tran Corp. An investment amount between 20% and 50% of the outstanding common stock of the investee is presumed to represent significant influence. The equity method is appropriate when the investor can exercise significant influence over the investee.